In an (x, y) coordinate system, write the equation of the vertical line passing
through the point of intersection of 3x + 4y = 1 and x + 3y = 7 .

Answer by nerdybill(7384)

In an (x, y) coordinate system, write the equation of the vertical line passing
through the point of intersection of 3x + 4y = 1 and x + 3y = 7 .
.
First, find out where the two lines intersect:
3x + 4y = 1
x + 3y = 7
.
Multiply second equation by -3:
3x + 4y = 1
-3x - 9y = -21
.
Add the two equations together:
3x + 4y = 1
-3x - 9y = -21
----------------
-5y = -20
y = 4
.
Use the second equation to find x:
x + 3y = 7
x + 3(4) = 7
x + 12 = 7
x = -5 (which is also the vertical line they are looking for)